Tohban Report 2015-09-09
Tohban Reports | |
---|---|
Start Date: | 2 Sep 2015 |
End Date: | 9 Sep 2015 |
Tohban: | Mitsuo Oka |
Tohban email: | moka@ssl.berkeley.edu |
Next Tohban: | TBD |
List all reports |
Solar Activity
Solar activity was extremely low with only 1 C-class flare and 9 B-class flare. There are currently 9 ARs on the disk but they are small and stable. The activity is expected to remain low.
How many GOES flares occurred?
Flares above B, C, M, X class were 9 1 0 0
And how many of these are listed in the RHESSI flare list?
Flares above B, C, M, X class were 3 0 0 0
And how many had EXCELLENT coverage?
Flares above B, C, M, X class were 3 0 0 0
There were RHESSI flares/GOES flares 8 / 10 over the time range 01-Sep-15 08-Sep-15
Memory Management
The SSR fill level has been decreasing gradually. It reached a maximum of ~46% and a minimum of about ~8% during the week.
Spacecraft Status
The cold plate 1 and 2 are at about 134.9 K and 133.3 K respectively.
Data Gaps
There was no significant data gaps.
Detector issues
All detectors remained stable almost through out the week, but the livetimes of D4 (front), D1 (rear) and D6 (rear) decreased rapidly on the last day of the week. The following changes were made:
D4 front-fast threshold raised from 0xF0 to 0xFF
D1 rear-fast threshold raised from 0x70 to 0x80
D6 rear-fast threshold raised from 0xE0 to 0xF0
D4 HV lowered from 4289 V to 3900 V
Other notes
Spacecraft Management
Three options were discussed regarding the attenuator issue:
Option 1: Leave att. in and pull out manually after exit of eclipse Option 2: Leave att. in and pull out manually every 15 min Option 3: Raise slow (fast) threshold for D7, trying to recover livetime and use D7 only for logic
A decision was made and Option 1 was selected.
Decimation | Active/Vigorous |
Night time data (fronts) | +- 4 minutes |
Night time data (rears) | +- 4 minutes |
Require extra passes? | No |
Requirement for moving pointer? | No |
Attenuator operation | Normal |